OpenStudy (anonymous):

simplify the rational exprression. 27x-21/7-9x

OpenStudy (anonymous):

start with \[\frac{3(9x-7)}{7-9x}\]

OpenStudy (anonymous):

and since \(\frac{9x-7}{7-9x}=-\frac{9x-7}{9x-7}=-1\) you are left with \(-3\)